# https://www.mmlab-ntu.com/project/edgesam

### v2

@@ 3.2.3 Granularity Priors
With the image encoder staying frozen, we build a lightweight region proposal network (RPN) [42] on top of it, which consists of a feature pyramid network (FPN) [29] and a shared detection head. For efficiency, we follow the design proposed by EfficientDet [47]. The RPN is trained on a specific dataset, e.g., COCO [31], to capture its granularity prior. During inference, we merge the proposal boxes whose centers are K nearest neighbors of the point prompts weighted by their confidence scores. Finally, we combine the merged box with the point input together as the prompt that inputs to the mask decoder.
@@ 3.3 Training and Application
Training Pipeline. We divide the training into three stages. In the first stage, we perform encoder-only knowledge distillation on 1% SA-1B images following MobileSAM [61]. In the second stage, we apply prompt-in-the-loop distillation on the same image set but with point and box prompts as part of the inputs. In the final stage, which is optional, we freeze modules other than the lightweight RPN and train with the class-agnostic ground-truth boxes with the commonly used focal loss [30] and Huber loss [22].
Inference and On-Device Demo. EdgeSAM is capable of initiating inference with either point or box prompts, and, akin to SAM, it can progressively incorporate additional points for further refinement. To show the practical utility of EdgeSAM, we have developed an on-device demonstration. We highly recommend exploring this demo to fully appreciate EdgeSAM’s capabilities.
@@ 4 Experiments
Accuracy. We divide the accuracy measurement into three scenarios according to the prompt type and source. (1) We use the ground-truth box as the initial prompt and iteratively add more point prompts in the wrongly segmented area for refinement. (2) Similar to (1) but using the center point of the ground-truth mask as the initial prompt. (3) An external object detector is leveraged to provide box prompts. For (1) and (2), we calculate the mIoU across all instances, and for (3), we report mAP and boundary IoU [9]. For all evaluations, we take the output of the first mask token as the prediction [^2] Since EfficientSAM only supports outputting multiple masks, we select the mask with the highest score as the final prediction for it..
Table 1: Efficiency comparison. FLOPs are calculated based on the $1024\times 1024$ input resolution. Numbers denoted by * are referred from MobileSAM [61]. Note that, MobileSAM and EfficientSAM-Ti share the same network architecture.
@@
Datasets. We consider SA-1B [24], COCO [31], and LVIS [12] to evaluate both in-domain performance and zero-shot transferability. For distillation, 1% images of the whole SA-1B dataset are used. We construct a test set, named SA-1K, from SA-1B. In particular, we randomly select 1,000 images from SA-1B, which are not used during training. For each image, we randomly sample 64 instances to avoid memory overflow. For COCO and LVIS, we use their validation sets.
@@ 4.1 Implementation Details
As mentioned in Sec. 3.3, we split the training into three stages. In the first stage, we train on 1% of SA-1B images with encoder-only distillation loss $L_{p}$. We train the model for 10 epochs with the batch size set to 64. We adopt the AdamW optimizer [33] and set the initial learning rate to 1.25e-2. We adopt the cosine decay schedule, which gradually decreases the learning rate to 6.25e-7. In the second stage, we load the encoder weights trained in stage one and inherit decoder weights from SAM. We then train with the prompt-in-the-loop distillation loss $L_{d}$ on 1% SA-1B for 5 epochs. The batch size is 16, and the learning rate decays from 1e-4 to 1e-5. We set the maximum allowed instances per image to 16, and the number of loops for prompt sampling is 1. In the final optional stage for RPN training, we follow the 1x schedule of MMDetection [6] on COCO [31].
@@ 4.2 Quantitative Results
Efficiency. In Tab. 1, we compare the FPS of SAM, FastSAM, MobileSAM, EfficientSAM-Ti, and EdgeSAM on both the desktop and mobile platforms. When evaluating with NVIDIA 2080 Ti, we compile the models with ONNX to avoid the overhead of the Python interpreter, and we make sure the GPU is running at 100% payload. On iPhone 14, we use the Core ML Tools [1] as the compiler and measure the throughput with the benchmark tools provided by Xcode. As shown in Tab. 1, the proposed EdgeSAM is significantly faster than other methods, especially on the mobile platform. In particular, EdgeSAM is over 37 times faster than SAM on NVIDIA 2080 Ti and over 7 times faster than MobileSAM and EfficientSAM on iPhone 14. From the benchmark report, we observe that MobileSAM and EfficientSAM are not well optimized by on-device AI accelerators, such as ANE, which explains the vast differences between desktop and mobile platforms. In addition, EfficientSAM requires a huge training cost including the 400 epochs on ImageNet-1K (IN) and 5 epochs on the entire SA-1B dataset.
@@
Prompting with Center Points. Similar to prompting with GT boxes, EdgeSAM shows a clear advantage over MobileSAM and EfficientSAM in most cases, as shown in Tab. 3. However, as we have discussed in Sec. 3.2.3, with ambiguous prompts, such as a single point, the original SAM does not always produce masks in the desired granularity. In addition, EdgeSAM is trained on the SA-1B dataset, whose granularity distribution is very different from that of COCO. Therefore, with the proposed task-aware distillation, EdgeSAM is more likely to fit the SAM granularity distribution when handling ambiguous prompts than MobileSAM. To explicitly capture such a granularity prior, we propose a lightweight RPN, which works as expected and makes the single point performance on COCO rise from 48.0 to 54.3. Besides, when compared with EfficientSAM, we would like to highlight that it is trained on a 100 times larger dataset than EdgeSAM and uses a larger backbone model. Later in the ablation studies (Sec. 4.3), we show that EdgeSAM can also benefit from larger train sets and models.
Table 5: Ablation studies. If not explicitly stated, we report the mIoUs on SA-1K and FPS on a single NVIDIA 2080 Ti. For more ablation studies, please refer to Appendix 0.B.
@@
(d) Generalizability of the proposed prompt-in-the-loop KD.
@@
(f) Frozen modules and LoRA [19]. F denotes freezing and L means applying LoRA during training.
Encoder Decoder SA-1K COCO
@@
(h) Train stages.
Box Center
Joint 81.2 63.2
2-stage 83.1 67.4
@@
w/ Structured [32] 81.6 65.2
w/ Channel-wise [44] 81.8 64.9
Prompting with External Object Detectors. We finally combine all the SAM variants with an external object detector (We use Detic [69] from MMDetection [6] and ViTDet-H [28] from Detectron2 [54]) and evaluate their mAP performance on COCO. This setting examines a practical use case where automatic segmentation is preferred over interactive segmentation, e.g., generating the mask within the tracking box. We use an object detection framework to generate bounding boxes and associated class labels, ensuring that the observed variations in mAP are attributed solely to the quality of the masks generated. This approach also evaluates the resilience of our model in handling imprecise bounding box inputs. We also report boundary IoU [9] to reflect the boundary accuracy. Tab. 4 shows that EdgeSAM outperforms MobileSAM by considerable margins and is on par with EfficientSAM with smaller training and inference costs. However, it still lags clearly behind SAM. This performance gap may reflect not only inherent limitations in model capacity but also a consequence of training exclusively with ground-truth boxes, which could lead to discrepancies during inference. Addressing these differences and exploring their implications is a promising direction for future research.
Figure 3: Qualitative results of EdgeSAM with point and box prompts. The green and red stars indicate the positive and negative points respectively. More comparative results between EdgeSAM and SAM are provided in Appendix 0.C.
4.3 Ablation Studies
Choice of the Backbone. As mentioned in Sec. 3.2.1, the output resolution of SAM and that of efficient backbones are misaligned. MobileSAM proposed to remove the downsampling operations in the last several stages, while we experiment with fusing low-resolution feature maps with higher ones through an FPN [29]. In addition, we compare the performance-speed trade-offs across efficient backbones following the ViT-based [53], CNN-based [50], and hybrid [3] designs. As shown in Tab. 5(c), the purely CNN-based RepViT-M1 with FPN for resolution alignment achieves the best balance. Moreover, as many on-device AI accelerators, such as ANE, are highly optimized for CNNs, the speed gap becomes larger when deployed on edge devices. For example, on iPhone 14, with FPN for resolution alignment, RepViT-M1 only takes 14 ms to encode a $1024\times 1024$ input, which is 14x and 4x faster than TinyViT-5M and EfficientViT-B1 respectively. Therefore, our EdgeSAM adopts it as the image encoder.
Generalizability of the Proposed Prompt-in-the-Loop KD. While we opt for RepViT given it is more optimized for current mobile platforms (empirically proved by the previous ablation), the proposed prompt-in-the-loop knowledge distillation method is not tied to a specific backbone. Therefore, in Tab. 5(d), we show that with the proposed KD method, the performance of the resulting model is consistently and significantly improved across TinyViT, EfficientViT, and RepViT without any additional cost during inference, which demonstrates the generalizability of our method.
Other Ablations. In Tab. 5(f), we test several configurations regarding module freezing and LoRA [19]. In particular, LoRA is applied to the query and value projection layers of the attention blocks in the decoder. Results show that freezing the decoder during distillation yields the best in-domain accuracy while fine-tuning all the modules generalizes better on other datasets. To benefit more practical settings, we opt to fine-tune both the encoder and decoder. Furthermore, we vary the number of prompt sampling loops during distillation, in Tab. 5(g), we find one additional loop is good enough, thus setting the loop number to be 1 by default. In addition, due to the optimal optimization settings being different for encoder-only distillation and prompt-in-the-loop distillation, we find joint-training yields sub-optimal results compared to training in two stages (see Tab. 5(h)). Therefore, we propose to train them sequentially. Finally, in Tab. 5(i), we explore several ways to improve the performance of encoder-only distillation, including training for a longer schedule and applying knowledge distillation losses designed for dense prediction tasks [32, 44]. While they boost the performance of point prompts, that of box prompts slightly drops, indicating an upper-bound performance of encoder-only distillation. This motivates us to explore prompt-aware distillation.
4.4 Qualitative Results
In Fig. 3, we visualize the mask quality of EdgeSAM with one point, two points, and one box as the prompts, respectively. Note that none of the images are from the SA-1B dataset that EdgeSAM is trained on, demonstrating its zero-shot generalization ability. We provide more visual comparisons between SAM and EdgeSAM in Appendix 0.C.
5 Conclusion and Discussion
In this paper, we present EdgeSAM, the first SAM variant that runs in real-time on edge devices. We achieve this by distilling SAM into a lightweight CNN-based architecture. Our preliminary experiments show that existing distillation schemes, that only involve the image encoder thus being task-agnostic, fail to reveal the full knowledge spectrum of SAM to the student model. Thus, we propose a prompt-in-the-loop knowledge distillation method that considers both the encoder and decoder of SAM and provides task-specific supervision signals. Experiments across SA-1B, COCO, and LVIS under various settings demonstrate the effectiveness of EdgeSAM in terms of efficiency and accuracy compared with SAM and other accelerated variants.
Despite being able to run EdgeSAM at a decent speed on edge devices, we still have several research directions, that could potentially provide boosts, left to be investigated, including quantization, model pruning, on-device optimization, mixed-precision inference, etc. Besides, we haven’t applied any augmentation during training, thus proper data or prompt augmentation might also be promising directions. We hope EdgeSAM will encourage more real-world applications with its on-device interactive segmentation capability and will continue working on further improvements.
Appendix 0.A Effectiveness of the Prompt-in-the-Loop Knowledge Distillation
In the main paper, we show the quantitative results of prompt-in-the-loop knowledge distillation (prompt-KD) compared with encoder-only knowledge distillation methods, e.g., MobileSAM. Here, we supplement the visual results to further analyze the effectiveness of prompt-KD. In specific, we compare EdgeSAM with MobileSAM with both points and boxes as prompts. Since EdgeSAM and MobileSAM adopt different backbones, we add a baseline that follows the same network architecture as EdgeSAM but is trained with MobileSAM’s encoder-only distillation method. As shown in Fig. 4, compared with task-agnostic KD methods, models trained with prompt-KD follow the user’s intent more precisely, e.g., MobileSAM and baseline fail to segment both persons on the bench. Besides, prompt-KD involves implicit hard example mining by iteratively sampling new prompts in the wrongly segmented area, which leads to tighter boundaries and more accurate results in difficult regions, e.g., the sculpture and bicycle in Fig. 4. These results align well with our assumption that task-agnostic encoder distillation fails to capture the full knowledge embodied in SAM and involving prompts in the loop during distillation helps capture the intricate dynamics between user input and mask generation.
Appendix 0.B More Ablation Studies
Training Datasets. In our default setting, we train EdgeSAM with 1% images of the SA-1B [24] dataset. In Tab. 6(a), we study the impact of training datasets on transferability. Specifically, we keep the model and training pipeline the same while switching the training dataset among 1% SA-1B, COCO [31], and LVIS [12]. Note that all datasets contain a similar number of images (approximately 110K). Results show that training with SA-1B archives the best performance across all test sets with box prompts, while with point prompts, training and testing in the same domain yield the best result. This suggests that with informative prompts, such as boxes, distillation on the SA-1B dataset is sufficient for zero-shot transfer. However, with ambiguous prompts, such as a single point, efforts must be made to capture the granularity priors of the testing datasets. This motivates us to leverage an additional RPN to capture such priors explicitly.
Table 6: More ablation studies. Box and center denote testing EdgeSAM with GT-box and center of the GT-mask as the prompt respectively. COCO mAP and boundary IoU [9] are obtained with the Detic [69] object detector providing box prompts and class labels (box mAP 47.4). Note that, for all the ablations, we set the number of prompt sampling loops to 0 (no re-sampling). SA-1B* means 1% SA-1B. Bold font indicates the best number and the default configuration we adopt.
(a) Training datasets.
Dataset SA-1K COCO LVIS
Box Center Box Center Box Center
SA-1B* 83.1 67.4 76.7 48.2 76.1 53.5
COCO 81.3 63.1 76.1 50.8 74.7 54.2
LVIS 82.4 63.6 75.7 50.2 75.4 54.9
(b) Decoder losses.
@@ Appendix 0.B More Ablation Studies
(c) Number of prompts. In this ablation, we freeze the mask decoder of EdgeSAM.
#Prompt SA-1K COCO
Box Center AP BIoU
8 83.6 67.6 34.3 21.7
16 83.5 67.7 34.6 21.9
32 83.6 67.6 34.3 21.6

### v3

@@ 3.2.3 Granularity Priors
With the image encoder staying frozen, we build a lightweight region proposal network (RPN) (Ren et al, 2015) on top of it, which consists of a feature pyramid network (FPN) (Lin et al, 2017a) and a shared detection head. For efficiency, we follow the design proposed by EfficientDet (Tan et al, 2020). The RPN is trained on a specific dataset, e.g., COCO (Lin et al, 2014), to capture its granularity prior. During inference, we merge the proposal boxes whose centers are K nearest neighbors of the point prompts weighted by their confidence scores. Finally, we combine the merged box with the point input together as the prompt that inputs to the mask decoder.
@@ 3.3 Training and Application
Training Pipeline. We divide the training into three stages. In the first stage, we perform encoder-only knowledge distillation on 1% SA-1B images following MobileSAM (Zhang et al, 2023a). In the second stage, we apply prompt-in-the-loop distillation on the same image set but with point and box prompts as part of the inputs. In the final stage, which is optional, we freeze modules other than the lightweight RPN and train with the class-agnostic ground-truth boxes with the commonly used focal loss (Lin et al, 2017b) and Huber loss (Huber, 1992).
Inference and On-Device Demo. EdgeSAM is capable of initiating inference with either point or box prompts, and, akin to SAM, it can progressively incorporate additional points for further refinement. To show the practical utility of EdgeSAM, in demo.mov, we provide a live demonstration of running EdgeSAM on an iPhone. We develop an iOS application and convert the trained EdgeSAM model to a Core ML model serving as the backend (we use the coremltools (cor, 2021) for this purpose). The demo video also showcases the qualitative results of applying EdgeSAM to challenging scenes including crowded tiny objects, night-time conditions, different domains, etc.. Our demonstration device is an iPhone 13 running on iOS 17.1.
@@ 4 Experiments
Accuracy. We divide the accuracy measurement into three scenarios according to the prompt type and source. (1) We use the ground-truth box as the initial prompt and iteratively add more point prompts in the wrongly segmented area for refinement. (2) Similar to (1) but using the center point of the ground-truth mask as the initial prompt. (3) An external object detector is leveraged to provide box prompts. For (1) and (2), we calculate the mIoU across all instances, and for (3), we report mAP and boundary IoU (Cheng et al, 2021). For all evaluations, we take the output of the first mask token as the prediction [^2] Since EfficientSAM only supports outputting multiple masks, we select the mask with the highest score as the final prediction for it..
Datasets. We consider SA-1B (Kirillov et al, 2023), COCO (Lin et al, 2014), and LVIS (Gupta et al, 2019) to evaluate both in-domain performance and zero-shot transferability. For distillation, 1% images of the whole SA-1B dataset are used. We construct a test set, named SA-1K, from SA-1B. In particular, we randomly select 1,000 images from SA-1B, which are not used during training. For each image, we randomly sample 64 instances to avoid memory overflow. For COCO and LVIS, we use their validation sets.
Table 1: Efficiency comparison. FLOPs are calculated based on the $1024\times 1024$ input resolution. Numbers denoted by * are referred from MobileSAM (Zhang et al, 2023a). Note that, MobileSAM and EfficientSAM-Ti share the same network architecture.
@@ 4.1 Implementation Details
As mentioned in Sec. 3.3, we split the training into three stages. In the first stage, we train on 1% of SA-1B images with encoder-only distillation loss $L_{p}$. We train the model for 10 epochs with the batch size set to 64. We adopt the AdamW optimizer (Loshchilov and Hutter, 2019) and set the initial learning rate to 1.25e-2. We adopt the cosine decay schedule, which gradually decreases the learning rate to 6.25e-7. In the second stage, we load the encoder weights trained in stage one and inherit decoder weights from SAM. We then train with the prompt-in-the-loop distillation loss $L_{d}$ on 1% SA-1B for 5 epochs. The batch size is 16, and the learning rate decays from 1e-4 to 1e-5. We set the maximum allowed instances per image to 16, and the number of loops for prompt sampling is 1. In the final optional stage for RPN training, we follow the 1x schedule of MMDetection (Chen et al, 2019) on COCO (Lin et al, 2014). Note that, for both the first and second stages, following TinyViT (Wu et al, 2022), we cache output embeddings from the teacher’s encoder to disk to minimize the training overhead. In addition, since EdgeSAM is data efficient and only requires 1% of SA-1B samples, the total training cost is lower than MobileSAM.
@@ 4.2 Quantitative Results
Efficiency. In Tab. 1, we compare the FPS of SAM, FastSAM, MobileSAM, EfficientSAM-Ti, and EdgeSAM on both the desktop and mobile platforms. When evaluating with NVIDIA 2080 Ti, we compile the models with ONNX to avoid the overhead of the Python interpreter, and we make sure the GPU is running at 100% payload. On iPhone 14, we use the Core ML Tools (cor, 2021) as the compiler and measure the throughput with the benchmark tools provided by Xcode. As shown in Tab. 1, the proposed EdgeSAM is significantly faster than other methods, especially on the mobile platform. In particular, EdgeSAM is over 37 times faster than SAM on NVIDIA 2080 Ti and over 7 times faster than MobileSAM and EfficientSAM on iPhone 14. From the benchmark report, we observe that MobileSAM and EfficientSAM are not well optimized by on-device AI accelerators, such as ANE, which explains the vast differences between desktop and mobile platforms. In addition, EfficientSAM requires a huge training cost including the 400 epochs on ImageNet-1K (IN) and 5 epochs on the entire SA-1B dataset.
@@
Prompting with Center Points. Similar to prompting with GT boxes, EdgeSAM shows a clear advantage over MobileSAM and EfficientSAM in most cases, as shown in Tab. 3. However, as we have discussed in Sec. 3.2.3, with ambiguous prompts, such as a single point, the original SAM does not always produce masks in the desired granularity. In addition, EdgeSAM is trained on the SA-1B dataset, whose granularity distribution is very different from that of COCO. Therefore, with the proposed task-aware distillation, EdgeSAM is more likely to fit the SAM granularity distribution when handling ambiguous prompts than MobileSAM. To explicitly capture such a granularity prior, we propose a lightweight RPN, which works as expected and makes the single point performance on COCO rise from 48.0 to 54.3. As we do not observe a granularity degradation on other datasets except for COCO, we refrain from applying RPN on them to keep the model simple and efficient. Besides, when compared with EfficientSAM, we would like to highlight that it is trained on a 100 times larger dataset than EdgeSAM and uses a larger backbone model. Later in the ablation studies (Sec. 4.3), we show that EdgeSAM can also benefit from larger train sets and models.
Table 5: Ablation studies. If not explicitly stated, we report the mIoUs on SA-1K and FPS on a single NVIDIA 2080 Ti. COCO mAP and boundary IoU (Cheng et al, 2021) are obtained with the Detic (Zhou et al, 2022) object detector providing box prompts and class labels (box mAP 47.4). Bold font indicates the best number and the default configuration we adopt.
@@
(d) Generalizability of prompt-in-the-loop KD.
@@
(f) Frozen modules and LoRA. F denotes freezing and L means applying LoRA during training.
Enc. Dec. SA-1K COCO
@@
(h) Number of prompts. Mask decoder is frozen.
#Prompt SA-1K COCO
Box Center AP BIoU
8 83.6 67.6 34.3 21.7
16 83.5 67.7 34.6 21.9
32 83.6 67.6 34.3 21.6
@@
w/ Structured 81.6 65.2
w/ Channel-wise 81.8 64.9
(j) Decoder losses.
@@
(k) Train stages.
Box Center
Joint 81.2 63.2
2-stage 83.1 67.4
(l) Training datasets. SA-1B* means 1% SA-1B.
Dataset SA-1K COCO LVIS
Box Center Box Center Box Center
SA-1B* 83.1 67.4 76.7 48.2 76.1 53.5
COCO 81.3 63.1 76.1 50.8 74.7 54.2
LVIS 82.4 63.6 75.7 50.2 75.4 54.9
Prompting with External Object Detectors. We finally combine all the SAM variants with an external object detector (We use Detic (Zhou et al, 2022) from MMDetection (Chen et al, 2019) and ViTDet-H (Li et al, 2022b) from Detectron2 (Wu et al, 2019)) and evaluate their mAP performance on COCO. This setting examines a practical use case where automatic segmentation is preferred over interactive segmentation, e.g., generating the mask within the tracking box. We use an object detection framework to generate bounding boxes and associated class labels, ensuring that the observed variations in mAP are attributed solely to the quality of the masks generated. This approach also evaluates the resilience of our model in handling imprecise bounding box inputs. We also report boundary IoU (Cheng et al, 2021) to reflect the boundary accuracy. Tab. 4 shows that EdgeSAM outperforms MobileSAM by considerable margins and is on par with EfficientSAM with smaller training and inference costs. However, it still lags clearly behind SAM. This performance gap may reflect not only inherent limitations in model capacity but also a consequence of training exclusively with ground-truth boxes, which could lead to discrepancies during inference. Addressing these differences and exploring their implications is a promising direction for future research.
4.3 Ablation Studies
Choice of the Backbone. As mentioned in Sec. 3.2.1, the output resolution of SAM and that of efficient backbones are misaligned. MobileSAM proposed to remove the downsampling operations in the last several stages, while we experiment with fusing low-resolution feature maps with higher ones through an FPN (Lin et al, 2017a). In addition, we compare the performance-speed trade-offs across efficient backbones following the ViT-based (Wu et al, 2022), CNN-based (Wang et al, 2023b), and hybrid (Cai et al, 2023) designs. As shown in Tab. 5(c), the purely CNN-based RepViT-M1 with FPN for resolution alignment achieves the best balance. Moreover, as many on-device AI accelerators, such as ANE, are highly optimized for CNNs, the speed gap becomes larger when deployed on edge devices. For example, on iPhone 14, with FPN for resolution alignment, RepViT-M1 only takes 14 ms to encode a $1024\times 1024$ input, which is 14x and 4x faster than TinyViT-5M and EfficientViT-B1 respectively. Therefore, our EdgeSAM adopts it as the image encoder.
Generalizability of the Proposed Prompt-in-the-Loop KD. While we opt for RepViT given it is more optimized for current mobile platforms (empirically proved by the previous ablation), the proposed prompt-in-the-loop knowledge distillation method is not tied to a specific backbone. Therefore, in Tab. 5(d), we show that with the proposed KD method, the performance of the resulting model is consistently and significantly improved across TinyViT, EfficientViT, and RepViT without any additional cost during inference, which demonstrates the generalizability of our method. Note that, RepViT-M1 without our prompt-in-the-loop KD is essentially RepViT-SAM (Wang et al, 2023a). Therefore, EdgeSAM outperforms RepViT-SAM by 1.0 and 2.9 mIoUs with box and center point as the prompt, respectively  [^3] The original RepViT-SAM adopts RepViT-M2.3 as the encoder while here we compare the performance with the RepViT-M1 backbone..
Frozen Modules and LoRA. In Tab. 5(f), we test several configurations regarding module freezing and LoRA (Hu et al, 2022). In particular, LoRA is applied to the query and value projection layers of the attention blocks in the decoder. Results show that freezing the decoder during distillation yields the best in-domain accuracy while fine-tuning all the modules generalizes better on other datasets. To benefit more practical settings, we opt to fine-tune both the encoder and decoder.
Number of Prompt Sampling Loops. In Tab. 5(g), we find that one additional loop achieves the best cost-performance trade-off, and we conjecture that with one correction, the student is already doing better enough to match the teacher, that is, the prediction difference between student and teacher does not leave enough sampling area to generate new prompts from. Thus we set the loop number to be 1 by default. Note that when setting the number of loops to 0, the decoder is still inferred once and gets trained, which is different from encoder-only KD where the decoder is not involved during training.
Number of Prompts. Given that each sample of SA-1B contains around 100 instances, training all of them in a single batch causes VRAM overflow. Therefore, we randomly sample a subset for each iteration. In Tab. 5(h), we study the impact of the number of prompts for sampling and find that 16 prompts achieve the best trade-off. Thus, we set it as the default.
Encoder-Only KD. In Tab. 5(i), we explore several ways to improve the performance of encoder-only distillation, including training for a longer schedule and applying knowledge distillation losses designed for dense prediction tasks (Liu et al, 2019; Shu et al, 2021). While they boost the performance of point prompts, that of box prompts slightly drops, indicating an upper-bound performance of encoder-only distillation. This motivates us to explore prompt-aware distillation.